# Are the Majority of Public Computational Notebooks Pathologically Non-Executable?

### v1

@@ III Research Questions
Similar to different notions of executability, we also introduce different degrees of executability. Instead of considering it as a binary metric, we measure executability on a continuous spectrum, defined as a ratio between the number of cells executed until the first error and the total number of cells in a notebook. 100% executability refers to fully executable notebook, whereas $<$100% refers to partially executable notebooks. A fully executable notebook does not guarantee reproducibility, i.e., a notebook must produce the same outputs as the original notebook. Reproducibility is beyond the scope of this work and is often not required for data-centric notebooks as they are expected to produce different outputs on different data. We explore the following research questions:
@@ V-C RQ3: Degree of Executability of Pathologically Non-executable Notebooks
>>>0-910-1920-2930-3940-4950-5960-6970-7980-8990-99002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,03810381{,}0381 , 038997997997997824824824824564564564564384384384384472472472472321321321321273273273273290290290290165165165165Executable Cells (%)# of Notebooks
@@
Our analysis reveals that, on average, 34.1% of the cells in pathologically non-executable notebooks are executable. For example, although the notebook [23] encounters “ValueError” during linear execution, it is still partially executable up to 10 cells or in 76.9% of total cells. Figure 7 shows the frequency of pathologically non-executable notebooks with different levels of partial executability. The Y-axis represents the number of notebooks, and the X-axis represents the percentage of cells successfully executed in those notebooks. For instance, we observe that nearly 17.2% of the pathologically non-executable notebooks (1,270) successfully execute more than half of the cells ($>$50%). On the other hand, 2,059 (27.9%) notebooks have 0% partial executability due to the pathological error in the first cell. Partial execution in nearly three-quarters of pathological non-executable notebooks demonstrates that even non-executable notebooks can be partially adopted, and dynamic analysis techniques can be applied to improve their executability and reproducibility further.
@@ V-D1 Addressing ModuleNotFound
Figure 8 shows the number of notebooks and the corresponding improvements in executability after applying this strategy. The Y-axis represents the number of restored notebooks, and the X-axis represents the increase in executable cells. The figure shows that 500 notebooks have $>$90% improvements in executability. For example, the notebook [24] initially encounters a “ModuleNotFound” error due to the absence of networkx module in the environment at cell 1. The environment requirement information is not provided in the repository. This approach restores the full executability of that notebook.
@@ V-D2 Addressing FileNotFound
$>$0-1011-2021-3031-4041-5051-6061-7071-8081-9091-10000200200200200400400400400600600600600800800800800847847847847584584584584298298298298298298298298264264264264138138138138193193193193220220220220260260260260500500500500Executable Cells (%)# of Notebooks
@@
$>$0-1011-2021-3031-4041-5051-6061-7071-8081-9091-10000100100100100200200200200300300300300400400400400392392392392327327327327158158158158122122122122767676765151515161616161565656565757575751515151Executable Cells (%)# of Notebooks
@@
$\geq$1000500-999300-499200-299150-199125-149100-12490-9980-8970-7960-6955-5950-5445-4940-4435-3930-3425-2920-2415-1910-144-90123.$10^{3}.$3073073073072,13521352{,}1352 , 1351941941941941141141141141031031031031,98019801{,}9801 , 9803313313313311,23912391{,}2391 , 2391,10011001{,}1001 , 10011311311311310910910910995959595142142142142266266266266121121121121106106106106169169169169120120120120128128128128949494942132132132131711711711712902902902901,85118511{,}8511 , 85116616616616691919191838383831,70917091{,}7091 , 709288288288288996996996996893893893893969696968585858567676767113113113113234234234234989898988585858514214214214210110110110110310310310372727272173173173173151151151151GitHub Stars # of Executable Notebooks Initial CountFinal Count
@@ VI Discussion
Pimentel et al.[3, 5] Empirical $>$1.4M N/A N/A N/A 76% Includes high-number of low-starred, rarely reused notebooks from GitHub
@@ VIII Conclusion
Computational notebooks, widely used for data science and ML/AI tasks, continually suffer from executability issues. Prior studies reporting the non-executability of notebooks rely on a rigid definition of executability, leading to over-estimating non-executable notebooks. For the first time, we introduce the notions of partial executability and pathological non-executability for notebooks, contextualizing executability to the notebook’s interactive computing paradigm. Our investigation finds 7,387 out of 42,546 notebooks are pathologically non-executable, while 27,272 can be restored given suitable execution environments. We leverage LLM-based error-driven restoration techniques to fully restored 5.4% and partially restored 12.8% of previously non-executable notebooks. These results offer key evidence that notebooks can benefit from LLM-based restoration, and partial executable notebooks are still valuable for broader code reuse practices.

### v3

@@ III Research Questions
Similar to different notions of executability, we also introduce different degrees of executability. Instead of considering it as a binary metric, we measure executability on a continuous spectrum, defined as a ratio between the number of cells executed until the first error and the total number of cells in a notebook. 100% executability refers to fully executable notebook, whereas <<100% refers to partially executable notebooks. A fully executable notebook does not guarantee reproducibility, i.e., a notebook must produce the same outputs as the original notebook. Reproducibility is beyond the scope of this work and is often not required for data-centric notebooks as they are expected to produce different outputs on different data. We explore the following research questions:
@@ V-C RQ3: Degree of Executability of Pathologically Non-executable Notebooks
>>0-910-1920-2930-3940-4950-5960-6970-7980-8990-9902002004004006006008008001,0001{,}0001,0381{,}038997997824824564564384384472472321321273273290290165165Executable Cells (%)# of Notebooks
@@
Our analysis reveals that, on average, 34.1% of the cells in pathologically non-executable notebooks are executable. For example, although the notebook [23] encounters “ValueError” during linear execution, it is still partially executable up to 10 cells or in 76.9% of total cells. Figure 7 shows the frequency of pathologically non-executable notebooks with different levels of partial executability. The Y-axis represents the number of notebooks, and the X-axis represents the percentage of cells successfully executed in those notebooks. For instance, we observe that nearly 17.2% of the pathologically non-executable notebooks (1,270) successfully execute more than half of the cells (>>50%). On the other hand, 2,059 (27.9%) notebooks have 0% partial executability due to the pathological error in the first cell. Partial execution in nearly three-quarters of pathological non-executable notebooks demonstrates that even non-executable notebooks can be partially adopted, and dynamic analysis techniques can be applied to improve their executability and reproducibility further.
@@ V-D1 Addressing ModuleNotFound
Figure 8 shows the number of notebooks and the corresponding improvements in executability after applying this strategy. The Y-axis represents the number of restored notebooks, and the X-axis represents the increase in executable cells. The figure shows that 500 notebooks have >>90% improvements in executability. For example, the notebook [24] initially encounters a “ModuleNotFound” error due to the absence of networkx module in the environment at cell 1. The environment requirement information is not provided in the repository. This approach restores the full executability of that notebook.
@@ V-D2 Addressing FileNotFound
>>0-1011-2021-3031-4041-5051-6061-7071-8081-9091-1000200200400400600600800800847847584584298298298298264264138138193193220220260260500500Executable Cells (%)# of Notebooks
@@
>>0-1011-2021-3031-4041-5051-6061-7071-8081-9091-1000100100200200300300400400392392327327158158122122767651516161565657575151Executable Cells (%)# of Notebooks
@@
[…]
@@ VI Discussion
Pimentel et al.[3, 5] Empirical >>1.4M N/A N/A N/A 76% Includes high-number of low-starred, rarely reused notebooks from GitHub
@@ VIII Conclusion
Computational notebooks, widely used for data science and ML/AI tasks, continually suffer from executability issues. Prior studies reporting the non-executability of notebooks rely on a rigid definition of executability, leading to overestimating non-executable notebooks. For the first time, we introduce the notions of partial executability and pathological non-executability for notebooks, contextualizing executability to the notebook’s interactive computing paradigm. Our investigation finds that 7,387 out of 42,546 notebooks are pathologically non-executable, while 27,272 can be restored given suitable execution environments. We leverage LLM-based error-driven restoration techniques to fully restore 5.4% and partially restore 12.8% of previously non-executable notebooks. These results offer key evidence that notebooks can benefit from LLM-based restoration, and partial executable notebooks are still valuable for broader code reuse practices.